# Beam Enumeration: Probabilistic Explainability For Sample Efficient Self-conditioned Molecular Design

### v1

@@ Abstract
Generative molecular design has moved from proof-of-concept to real-world applicability, as marked by the surge in very recent papers reporting experimental validation. Key challenges in explainability and sample efficiency present opportunities to enhance generative design to directly optimize expensive high-fidelity oracles and provide actionable insights to domain experts. Here, we propose Beam Enumeration to exhaustively enumerate the most probable sub-sequences from language-based molecular generative models and show that molecular substructures can be extracted. When coupled with reinforcement learning, extracted substructures become meaningful, providing a source of explainability and improving sample efficiency through self-conditioned generation. Beam Enumeration is generally applicable to any language-based molecular generative model and notably further improves the performance of the recently reported Augmented Memory algorithm, which achieved the new state-of-the-art on the Practical Molecular Optimization benchmark for sample efficiency. The combined algorithm generates more high reward molecules and faster, given a fixed oracle budget. Beam Enumeration is the first method to jointly address explainability and sample efficiency for molecular design. The code is available at https://github.com/schwallergroup/augmented_memory.
@@ 1 Introduction
Molecular discovery requires identifying candidate molecules possessing desired properties amidst an enormous chemical spaceSanchez-Lengeling and Aspuru-Guzik (2018). Generative molecular design has become a popular paradigm in drug discovery, offering the potential to navigate chemical space more efficiently with promise for accelerated discovery. Very recently, efforts have come to fruition and a large number of works have reported experimental validation of generated inhibitors, notably for both distribution learningMerk et al. (2018); Moret et al. (2021); Grisoni et al. (2021); Yu et al. (2021); Eguida et al. (2022); Li et al. (2022); Tan et al. (2021); Jang et al. (2022); Chen et al. (2022); Hua et al. (2022); Song et al. (2023); Moret et al. (2023); Ballarotto et al. (2023) and goal-directed generationKorshunova et al. (2022); Yoshimori et al. (2021); Zhavoronkov et al. (2019); Ren et al. (2023); Li et al. (2023); Salas-Estrada et al. (2023) approaches. Perhaps now more than ever, existing challenges in explainability and sample efficiency offer an avenue to propel generative molecular design towards outcomes that are not yet possible. Specifically, if one can elucidate why certain substructures or molecules satisfy a target objective, the model’s knowledge can be made actionable, for example, in an interplay with domain experts. Moreover, sample efficiency concerns with how many experiments, i.e., oracle calls, are required for a model to optimize the target objective. This is a pressing problem as the most informative high-fidelity oracles are computationally expensive, e.g., molecular dynamics (MD) for binding energy predictionWang et al. (2015); Moore et al. (2023). If a generative model can directly optimize these expensive oracles, the capabilities of generative design can be vastly advanced.
In this work, we propose Beam Enumeration to exhaustively enumerate the most probable token sub-sequences in language-based molecular generative models and show that valid molecular substructures can be extracted from these partial trajectories. We demonstrate that the extracted substructures are informative when coupled with reinforcement learning (RL) and show that this information can be made actionable to self-condition the model’s generation by only evaluating sampled molecules containing these substructures with the oracle. The results show significantly enhanced sample efficiency with an expected small trade-off in diversity. Beam Enumeration is the first method to jointly address explainability and sample efficiency. Our contribution is as follows:
@@
    We combine Beam Enumeration with the recently reported Augmented MemoryGuo and Schwaller (2023) optimization algorithm and show that the sample efficiency becomes sufficient (up to a 29-fold increase on the most challenging task) to find high reward molecules that satisfy a docking objective with only 2,000 oracle calls in three drug discovery case studies.
@@ 2 Related Work
Sample Efficiency in Molecular Design. Tailored molecular generation is vital for practical applications as every use case requires optimizing for a bespoke property profile. Over the past several years, so-called goal-directed generation has been achieved using a variety of architectures, including Simplified molecular-input line-entry system (SMILES)Weininger (1988)-based recurrent neural networks (RNNs)Olivecrona et al. (2017); Popova et al. (2018); Segler et al. (2018); Goel et al. (2021), generative adversarial networks (GANs)Goodfellow et al. (2014); Sanchez-Lengeling et al. (2017); Guimaraes et al. (2018), variational autoencoders (VAEs)Kingma and Welling (2022); Gómez-Bombarelli et al. (2018); Zhavoronkov et al. (2019), graph-based models You et al. (2019); Jin et al. (2020); Mercado et al. (2021); Atance et al. (2022), GFlowNetsBengio et al. (2021a), and genetic algorithmsFu et al. (2022). However, while all methods can be successful in optimizing for various properties, the oracle budget, i.e., how many oracle calls (computational calculations) were required to do so, is rarely reported. To address this, Gao et al. Gao et al. (2022) proposed the Practical Molecular Optimization (PMO)Gao et al. (2022) benchmark, which assesses 25 models across 23 tasks and enforces a budget of 10,000 oracle calls. Recently, Guo et al. proposed Augmented MemoryGuo and Schwaller (2023), which uses a language-based molecular generative model and achieves the new state-of-the-art on the PMO benchmark. In this work, Beam Enumeration is proposed as an addition to language-based molecular generative models, and we show that coupling it with Augmented Memory drastically improves the sample efficiency.
@@
Explainability for Molecules. Explainable AI (XAI)Gohel et al. (2021) to interpret and explain model predictions is a vital component for decision-making. Existing methods include Gradient-weighted Class Activation Mapping (Grad-CAM)Selvaraju et al. (2017), which uses gradient-based heat maps for convolutional layers and Local Interpretable Model-agnostic Explanations (LIME)Ribeiro et al. (2016), which uses a locally interpretable model. Other methods include permutation importanceAltmann et al. (2010), which measures the performance change when shuffling feature values, and SHAP valuesShapley (1953), which measure the average contribution of each feature to the model’s prediction. For molecules, the Molecular Model Agnostic Counterfactual Explanations (MMACE)Wellawatte et al. (2022) method was proposed to search for the most similar counterfactual (model predicts the opposite label) molecule. Recently, the pBRICSVangala et al. (2023) algorithm was proposed to decompose a molecule into functional groups and then applying Grad-CAM to explain matched molecular pairs (MMPs), i.e., pairs of molecules differing by a single chemical group. In this work, we aim to leverage explainability to achieve self-conditioned molecular generation and enhance sample efficiency. While existing XAI methods can work well provided a dataset, making the explanations actionable during a generative design experiment that relies on an interplay between chemical space exploration and oracle feedback is difficult.
@@
To address this limitation, we introduce Beam Enumeration, which extracts molecular substructures directly from the model’s token sampling probabilities and derives explainability from a generative probabilistic perspective that is modulated by reward feedback. Our approach is based on the fact that during a successful optimization trajectory, it must become increasingly likely to generate desirable molecules. It is thus reasonable to hypothesize that the most probable substructures are on track to receiving high reward. We verify this statement in the Results section.
@@ 3 Proposed Method: Beam Enumeration
In this section, each component of Beam Enumeration (Fig. 1) is described: the base molecular generative model, the Beam Enumeration algorithm, and how Beam Enumeration harnesses the model’s built-in explainability which can be used to improve sample efficiency through self-conditioned generation (further details on Beam Enumeration are presented in Appendix A).
Autoregressive Language-based Molecular Generative Model. The starting point of Beam Enumeration is any autoregressive language-based molecular generative model. The specific model used in this work is Augmented MemoryGuo and Schwaller (2023) which recently achieved the new state-of-the-art performance on the PMOGao et al. (2022) benchmark for sample efficiency, outperforming modern graph neural network-based approachesFu et al. (2020); Xie et al. (2021) and GFlowNetsBengio et al. (2021b). Augmented Memory builds on REINVENTOlivecrona et al. (2017); Blaschke et al. (2020a) which is a SMILES-basedWeininger (1988) RNN using long-short-term memory (LSTM) cellsHochreiter and Schmidhuber (1997). The optimization process is cast as an on-policy RL problem. We define the state space, $S_{t}$, as all intermediate token sequences and the action space, $A_{t}(s_{t})$, as the token sampling probabilities (conditioned on a given a state). $A_{t}(s_{t})$ is given by the policy, $\pi_{\theta}$, which is parameterized by the RNN. The objective is to iteratively update the policy such that token sampling, $A_{t}(s_{t})$, yields trajectories (SMILES) with increasing reward. Formally, sampling a SMILES, $x$, is given by the product of conditional state probabilities (Equation 1), and the token sampling is Markovian:
@@
Goal-directed generation is achieved by defining the Augmented Likelihood (Equation 2) where the Prior is the pre-trained model and $S$ is the objective function which yields a reward given a SMILES, $x$.
@@
Minimizing $L(\theta)$ is equivalent to maximizing the expected reward as shown previouslyGuo and Schwaller (2023); Fialková et al. (2022).
@@
Beam Enumeration. Beam Enumeration is proposed based on two facts: firstly, on a successful optimization trajectory, the model’s weights must change such that it becomes increasingly likely to generate high reward molecules. Secondly, generation involves sampling from conditional probability distributions. It is therefore reasonable to assume that the highest probability trajectories are more likely to yield high reward. Correspondingly, Beam Enumeration (Fig. 1) enumerates the top $k$ tokens (by probability) sequentially for $N$ beam steps (as it is infeasible to sample the full trajectories due to combinatorial explosion), resulting in an exhaustive set of token sub-sequences. We show that meaningful molecular substructures can be extracted from these sub-sequences, which we harness and demonstrate how it can be made actionable. We note the closest work to ours is the application of Beam SearchGraves (2012); Boulanger-Lewandowski et al. (2013) for molecular designMoret et al. (2021) to find the highest probability trajectories. Our work differs as the objective is not to find a small set of the most probable sequences. Rather, we exhaustively enumerate the highest probability sub-sequences to extract molecular substructures for self-conditioned generation.
@@
Probabilistic Explainability. Here, we describe how probabilistic explainability can be extracted from the exhaustive set of token sub-sequences. Usually, token sequences are only translated into SMILES once the sequence is complete, i.e., the “end” token has been sampled. We hypothesized that molecular substructures can be extracted from a given sub-sequence by iteratively considering every (sub)-sub-sequence (Fig. 1). For example, given the sub-sequence “ABCD”, the set of (sub)-sub-sequences are: “A”, “AB”, “ABC”, and “ABCD”. In practice, we only consider (sub)-sub-sequences with at least three characters (“ABC” and “ABCD”) since each character loosely maps to one atom and three is approximately the minimum for meaningful functional groups, e.g., “C=O”, a carbonyl. It is expected that not every sub-sequence possesses (sub)-sub-sequences mapping to valid molecular substructures. Still, we show that a sufficient signal can be extracted (Appendix C). Finally, we implement two types of substructures: Scaffold, which extracts the Bemis-MurckoBemis and Murcko (1996) scaffold and Structure, which extracts any valid substructure. In the Results section, we discuss the predictable difference in behavior.
@@
Self-conditioned Generation. The sub-sequences were enumerated by taking the most probable $k$ tokens, and the model’s weights should be updated such that high reward molecules are increasingly likely to be generated. Correspondingly, it is reasonable to posit that the most frequent molecular substructures are on track to becoming high reward full molecules and that the substructures themselves possess properties aligned with the target objective. The generative process can be self-conditioned to filter sampled batches for the presence of these molecular substructures and discard those that do not (Fig. 1).
@@ 4 Metrics
$Generative\;Yield=\sum_{g=1}^{G}\mathbb{I}[R(g)>T]$ (4)
Oracle Burden (Equation 5) is defined as the number of oracle calls ($c$) required to generate $N$ unique molecules above a reward threshold. This is a direct measure of sample efficiency as high reward molecules satisfy the target objective, and the metric becomes increasingly important with expensive high-fidelity oracles. In this work, all Oracle Burden metrics are computed by not allowing more than 10 molecules to possess the same Bemis-MurckoBemis and Murcko (1996) scaffold, thus also explicitly considering diversity in the generated set.
$Oracle\;Burden=c\mid\sum_{g=1}^{G}\mathbb{I}[R(g)>T]=N$ (5)
@@ 5 Results and Discussion
Figure 2: Illustrative experiment with the following multi-parameter optimization objective: maximize tPSA, molecular weight < 350 Da, number of rings $\geq$ 2. a. Augmented MemoryGuo and Schwaller (2023) reward trajectory with annotated top-4 (excluding benzene) most frequent molecular substructure scaffolds at varying epochs using Beam Enumeration. b. Examples of molecules with high reward.
@@
We first design an illustrative experiment to demonstrate the feasibility of Beam Enumeration to extract meaningful substructures and, in turn, enable self-conditioned generation. Next, three drug discovery case studies to design inhibitors against dopamine type 2 receptor (DRD2)Wang et al. (2018), MK2 kinaseArgiriadi et al. (2010), and acetylcholinesterase (AChE)Kryger et al. (1999) were performed to demonstrate real-world application. The key result we convey is that Beam Enumeration can be added directly to existing algorithms, and it both provides structural insights into why certain molecules receive high rewards and improves sample efficiency to not only generate more high reward molecules, but also faster, given a fixed oracle budget.

### v2

@@ Abstract
Generative molecular design has moved from proof-of-concept to real-world applicability, as marked by the surge in very recent papers reporting experimental validation. Key challenges in explainability and sample efficiency present opportunities to enhance generative design to directly optimize expensive high-fidelity oracles and provide actionable insights to domain experts. Here, we propose Beam Enumeration to exhaustively enumerate the most probable sub-sequences from language-based molecular generative models and show that molecular substructures can be extracted. When coupled with reinforcement learning, extracted substructures become meaningful, providing a source of explainability and improving sample efficiency through self-conditioned generation. Beam Enumeration is generally applicable to any language-based molecular generative model and notably further improves the performance of the recently reported Augmented Memory algorithm, which achieved the new state-of-the-art on the Practical Molecular Optimization benchmark for sample efficiency. The combined algorithm generates more high reward molecules and faster, given a fixed oracle budget. Beam Enumeration shows that improvements to explainability and sample efficiency for molecular design can be made synergistic. The code is available at https://github.com/schwallergroup/augmented_memory.
@@ 1 Introduction
Molecular discovery requires identifying candidate molecules possessing desired properties amidst an enormous chemical space (Sanchez-Lengeling & Aspuru-Guzik (2018). Generative molecular design has become a popular paradigm in drug discovery, offering the potential to navigate chemical space more efficiently with promise for accelerated discovery. Very recently, efforts have come to fruition and a large number of works have reported experimental validation of generated inhibitors, notably for both distribution learning (Merk et al. (2018); Moret et al. (2021); Grisoni et al. (2021); Yu et al. (2021); Eguida et al. (2022); Li et al. (2022); Tan et al. (2021); Jang et al. (2022); Chen et al. (2022); Hua et al. (2022); Song et al. (2023); Moret et al. (2023); Ballarotto et al. (2023) and goal-directed generation (Korshunova et al. (2022); Yoshimori et al. (2021); Zhavoronkov et al. (2019); Ren et al. (2023); Li et al. (2023); Salas-Estrada et al. (2023) approaches. Perhaps now more than ever, existing challenges in explainability and sample efficiency offer an avenue to propel generative molecular design towards outcomes that are not yet possible. Specifically, if one can elucidate why certain substructures or molecules satisfy a target objective, the model’s knowledge can be made actionable, for example, in an interplay with domain experts. Moreover, sample efficiency concerns with how many experiments, i.e., oracle calls, are required for a model to optimize the target objective. This is a pressing problem as the most informative high-fidelity oracles are computationally expensive, e.g., molecular dynamics (MD) for binding energy prediction (Wang et al. (2015); Moore et al. (2023). If a generative model can directly optimize these expensive oracles, the capabilities of generative design can be vastly advanced.
In this work, we propose Beam Enumeration to exhaustively enumerate the most probable token sub-sequences in language-based molecular generative models and show that valid molecular substructures can be extracted from these partial trajectories. We demonstrate that the extracted substructures are informative when coupled with reinforcement learning (RL) and show that this information can be made actionable to self-condition the model’s generation by only evaluating sampled molecules containing these substructures with the oracle. The results show significantly enhanced sample efficiency with an expected small trade-off in diversity. Beam Enumeration jointly addresses explainability and sample efficiency. Our contribution is as follows:
@@
    We combine Beam Enumeration with the recently reported Augmented Memory (Guo & Schwaller (2023) optimization algorithm and show that the sample efficiency becomes sufficient (up to a 29-fold increase) to find high reward molecules satisfying a docking objective with only 2,000 oracle calls in three drug discovery case studies.
@@ 2 Related Work
Sample Efficiency in Molecular Design. Tailored molecular generation is vital for practical applications as every use case requires optimizing for a bespoke property profile. Over the past several years, so-called goal-directed generation has been achieved using a variety of architectures, including Simplified molecular-input line-entry system (SMILES) (Weininger (1988)-based recurrent neural networks (RNNs) (Olivecrona et al. (2017); Popova et al. (2018); Segler et al. (2018); Goel et al. (2021), generative adversarial networks (GANs) (Goodfellow et al. (2014); Sanchez-Lengeling et al. (2017); Guimaraes et al. (2018), variational autoencoders (VAEs) (Kingma & Welling (2022); Gómez-Bombarelli et al. (2018); Zhavoronkov et al. (2019), graph-based models  (You et al. (2019); Jin et al. (2020); Mercado et al. (2021c); Atance et al. (2022), GFlowNets (Bengio et al. (2021a), and genetic algorithms (Fu et al. (2022a). However, while all methods can be successful in optimizing for various properties, the oracle budget, i.e., how many oracle calls (computational calculations) were required to do so, is rarely reported. To address this, Gao et al.  (Gao et al. (2022) proposed the Practical Molecular Optimization (PMO) benchmark, which assesses 25 models across 23 tasks and enforces a budget of 10,000 oracle calls. Recently, Guo et al. proposed Augmented Memory (Guo & Schwaller (2023), which uses a language-based molecular generative model and achieves the new state-of-the-art on the PMO benchmark. In this work, Beam Enumeration is proposed as an addition to language-based molecular generative models, and we show that coupling it with Augmented Memory drastically improves the sample efficiency.
@@
Explainability for Molecules. Explainable AI (XAI) (Gohel et al. (2021) to interpret and explain model predictions is a vital component for decision-making. Existing methods include Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al. (2017), which uses gradient-based heat maps for convolutional layers and Local Interpretable Model-agnostic Explanations (LIME) (Ribeiro et al. (2016), which uses a locally interpretable model. Other methods include permutation importance (Altmann et al. (2010), which measures the performance change when shuffling feature values, and SHAP values (Shapley (1953), which measure the average contribution of each feature to the model’s prediction. For molecules, the Molecular Model Agnostic Counterfactual Explanations (MMACE) (Wellawatte et al. (2022) method was proposed to search for the most similar counterfactual (model predicts the opposite label) molecule. Recently, the pBRICS (Vangala et al. (2023) algorithm was proposed to decompose a molecule into functional groups and then applying Grad-CAM to explain matched molecular pairs (MMPs), i.e., pairs of molecules differing by a single chemical group. In the context of generative models, previous works have explicitly addressed explainability (Guo et al. (2022); Fu et al. (2022b) and jointly with sample efficiency (Fu et al. (2022b). In this work, we aim to make explainability actionable during a generative design experiment.
@@
To achieve this, we introduce Beam Enumeration, which extracts molecular substructures directly from the model’s token sampling probabilities and derives explainability from a generative probabilistic perspective that is modulated by reward feedback. Our approach is based on the fact that during a successful optimization trajectory, it must become increasingly likely to generate desirable molecules. It is thus reasonable to hypothesize that the most probable substructures are on track to receiving high reward. We verify this statement in the Results section and show that Beam Enumeration can also jointly address explainability and sample efficiency.
@@ 3 Proposed Method: Beam Enumeration
In this section, each component of Beam Enumeration (Fig. 1) is described: the base molecular generative model, the Beam Enumeration algorithm, and how Beam Enumeration harnesses the model’s built-in explainability which can be used to improve sample efficiency through self-conditioned generation (further details on Beam Enumeration are presented in Appendix B).
Autoregressive Language-based Molecular Generative Model. The starting point of Beam Enumeration is any autoregressive language-based molecular generative model. The specific model used in this work is Augmented Memory (Guo & Schwaller (2023) which recently achieved the new state-of-the-art performance on the PMO (Gao et al. (2022) benchmark for sample efficiency, outperforming modern graph neural network-based approaches (Fu et al. (2020); Xie et al. (2021) and GFlowNets (Bengio et al. (2021b). Augmented Memory builds on REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020) which is a SMILES-based (Weininger (1988) RNN using long-short-term memory (LSTM) cells (Hochreiter & Schmidhuber (1997). The optimization process is cast as an on-policy RL problem. We define the state space, $S_{t}$, as all intermediate token sequences and the action space, $A_{t}(s_{t})$, as the token sampling probabilities (conditioned on a given state). $A_{t}(s_{t})$ is given by the policy, $\pi_{\theta}$, which is parameterized by the RNN. The objective is to iteratively update the policy such that token sampling, $A_{t}(s_{t})$, yields trajectories (SMILES) with increasing reward. Formally, sampling a SMILES, $x$, is given by the product of conditional state probabilities (Equation 1), and the token sampling is Markovian:
@@
The Augmented Likelihood is defined, (Equation 2) where the Prior is the pre-trained model and $S$ is the objective function which yields a reward given a SMILES, $x$.
@@
Minimizing $L(\theta)$ is equivalent to maximizing the expected reward as shown previously (Guo & Schwaller (2023); Fialková et al. (2022).
@@
Beam Enumeration. Beam Enumeration is proposed based on two facts: firstly, on a successful optimization trajectory, the model’s weights must change such that it becomes increasingly likely to generate high reward molecules. Secondly, generation involves sampling from conditional probability distributions. It is therefore reasonable to assume that the highest probability trajectories are more likely to yield high reward. Correspondingly, Beam Enumeration (Fig. 1) enumerates the top $k$ tokens (by probability) sequentially for $N$ beam steps (as it is infeasible to sample the full trajectories due to combinatorial explosion), resulting in an exhaustive set of token sub-sequences. We show that meaningful molecular substructures can be extracted from these sub-sequences, which we harness and demonstrate how it can be made actionable. We note the closest work to ours is the application of Beam Search (Graves (2012); Boulanger-Lewandowski et al. (2013) for molecular design (Moret et al. (2021) to find the highest probability trajectories. Our work differs as the objective is not to find a small set of the most probable sequences. Rather, we exhaustively enumerate the highest probability sub-sequences to extract molecular substructures for self-conditioned generation. We further detail the differences to Beam Search in Appendix H.
@@
Probabilistic Explainability. Here, we describe how probabilistic explainability can be extracted from the exhaustive set of token sub-sequences. Usually, token sequences are only translated into SMILES once the sequence is complete, i.e., the “end” token has been sampled. We hypothesized that molecular substructures can be extracted from a given sub-sequence by iteratively considering every (sub)-sub-sequence (Fig. 1). For example, given the sub-sequence “ABCD”, the set of (sub)-sub-sequences are: “A”, “AB”, “ABC”, and “ABCD”. In practice, we only consider (sub)-sub-sequences with at least three characters (“ABC” and “ABCD”) since each character loosely maps to one atom and three is approximately the minimum for meaningful functional groups, e.g., “C=O”, a carbonyl. It is expected that not every sub-sequence possesses (sub)-sub-sequences mapping to valid molecular substructures. Still, we show that a sufficient signal can be extracted (Appendix D). Finally, we implement two types of substructures: Scaffold, which extracts the Bemis-Murcko (Bemis & Murcko (1996) scaffold and Structure, which extracts any valid substructure. In the Results section, we discuss the predictable difference in behavior.
@@
Self-conditioned Generation. The sub-sequences were enumerated by taking the most probable $k$ tokens, and the model’s weights should be updated such that high reward molecules are increasingly likely to be generated. Correspondingly, it is reasonable to posit that the most frequent molecular substructures are on track to becoming high reward full molecules and that the substructures themselves possess properties aligned with the target objective. Beam Enumeration saves a Pool of these substructures and filters future sampled batches of molecules to contain them, discarding those that do not. Effectively, the generative process is self-conditioned as the model will only be updated by generated molecules containing the extracted substructures (Fig. 1).
@@ 4 Metrics
$Generative\;Yield=\sum_{g=1}^{G}\mathbb{I}[R(g)\textgreater T]$ (4)
Oracle Burden (Equation 5) is defined as the number of oracle calls ($c$) required to generate $N$ unique molecules above a reward threshold. This is a direct measure of sample efficiency as high reward molecules satisfy the target objective, and the metric becomes increasingly important with expensive high-fidelity oracles. In this work, all Oracle Burden metrics are computed by not allowing more than 10 molecules to possess the same Bemis-Murcko (Bemis & Murcko (1996) scaffold, thus also explicitly considering diversity in the generated set.
$Oracle\;Burden=c\mid\sum_{g=1}^{G}\mathbb{I}[R(g)\textgreater T]=N$ (5)
@@ 5 Results and Discussion
Figure 2: Illustrative experiment with the following multi-parameter optimization objective: maximize tPSA, molecular weight <350 Da, number of rings $\geq$ 2. a. Augmented Memory (Guo & Schwaller (2023) reward trajectory with annotated top-4 (excluding benzene) most frequent molecular substructure scaffolds at varying epochs using Beam Enumeration. b. Examples of molecules with high reward.
@@
We first design an illustrative experiment to demonstrate the feasibility of Beam Enumeration to extract meaningful substructures and, in turn, enable self-conditioned generation. Next, three drug discovery case studies to design inhibitors against dopamine type 2 receptor (DRD2) (Wang et al. (2018), MK2 kinase (Argiriadi et al. (2010), and acetylcholinesterase (AChE) (Kryger et al. (1999) were performed to demonstrate real-world application. The key result we convey is that Beam Enumeration can be added directly to existing algorithms, and it both provides structural insights into why certain molecules receive high rewards and improves sample efficiency to not only generate more high reward molecules, but also faster, given a fixed oracle budget.